OpenStudy (anonymous):

The manufacturing cost of a cell phone was $136. Sellers first marked up the cost by 25 percent. Sales were poor, so sellers decreased the markup to 17 percent. Calculate the old and new selling prices.

OpenStudy (anonymous):

old selling price = 136 + (136*25)/100 new selling price = 136 + (136*17)/100

OpenStudy (anonymous):

old price=170 new price=159